OpenStudy (anonymous):

if f(x) = 2x + 3 and g(x) = x^2, then g[f(-3)] = ?

OpenStudy (pottersheep):

Ok, first solve what f(-3) = 2x + 3 would be. Plug in -3 wherever there is an x

OpenStudy (pottersheep):

f(-3) = 2x + 3 = 2(-3) + 3 = -6 +3 = -3

OpenStudy (pottersheep):

Therefore, g[f(-3)] is equal to g (-3) (because f(-3) = -3)

OpenStudy (pottersheep):

g(-3) = x^2 = (-3)^2 = 9

OpenStudy (anonymous):

thank you

OpenStudy (pottersheep):

np :)